OpenStudy (anonymous):

The half-life of a radioactive isotope is 20.0 minutes. What is the total amount of a 1.00-gram sample of this sample of this isotope remaining after 1.00 hour?

OpenStudy (xishem):

Well, looking at it non-analytically... After 1.00 hours, how many half-lifes have passed? 3. So the amount of substance would have been halved three times... \[1.00g \rightarrow 0.50g \rightarrow 0.25g \rightarrow 0.125g\]

OpenStudy (xishem):

And from a more mathematical standpoint...\[N(t)=N_0\frac{1}{2}^{\frac{t}{t_{1/2}}}\]Where... N(t) is the amount of substance remaining after a time, t N_0 is the intial amount of substance t is the time that has surpassed t_{1/2} is the half-life of the substance Plugging some numbers in...\[N(1.00hr)=1.00g*\frac{1}{2}^\frac{60\min}{20.0\min}\]\[N(1.00hr)=0.125g\]

OpenStudy (anonymous):

Thank You soooooo much!!!